OpenStudy (anonymous):

find the domain and range. g(x)=2/x-1.

OpenStudy (anonymous):

Would the domain be all real numbers, but x cannot equal 1

OpenStudy (ash2326):

yeah, correct

OpenStudy (anonymous):

I just ned help with range

OpenStudy (ash2326):

okay, let me think just a moment

OpenStudy (anonymous):

ok:)

OpenStudy (anonymous):

May I help??

OpenStudy (ash2326):

feel free to join in always @waterineyes

OpenStudy (anonymous):

For finding range we have to find x in terms of g(x)..

OpenStudy (anonymous):

hmm

OpenStudy (ash2326):

oops I made a mistake

OpenStudy (anonymous):

Put g(x) = y for simplicity: \[(x-1)y = 2 \implies xy - y = 2 \implies xy = 2+y \implies x = \frac{2+y}{y}\] Now you know what y cannot take..

OpenStudy (anonymous):

Have you got or not??

OpenStudy (anonymous):

y cannot be 2?

OpenStudy (anonymous):

No dear check for denominator..

OpenStudy (anonymous):

Im not sure,,

OpenStudy (anonymous):

It is just like as you have found Domain.. y cannot be 0. Right or not??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so my range is just y can't be o

OpenStudy (anonymous):

0

OpenStudy (ash2326):

@waterineyes good work

OpenStudy (anonymous):

Yes all real except 0..

OpenStudy (anonymous):

thanks so much . I'm posting another one now so if you don't mind, plz look! thanks!

OpenStudy (anonymous):

Remember: For Domain, we find x.. For range, we find y or f(x)..

OpenStudy (anonymous):

@ash2326 just learning it from you..